OpenStudy (anonymous):

Find the zeros and multiplicity of y=(x+5)(x-2)^2

OpenStudy (anonymous):

y=(x+5)(x-2)^2 y = (x² - 3)(x² - 2) y = (x - 3)(x + 2)(x - 2)(x + 2) zeros: 3 and 2 multiplicity: of 2

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

@bibby @unknown101 @jigglypuff314

OpenStudy (anonymous):

http://prntscr.com/61hj1c

OpenStudy (anonymous):

yes I think you are correct :D wat grade is this stuff?

OpenStudy (anonymous):

Algebra 2, 11th grade.

OpenStudy (anonymous):

ko im in da same grade lol im 17 and the same class but I think its right

OpenStudy (anonymous):

Oh, are you onlined schooled?

OpenStudy (anonymous):

yup i go to k12 hbu?

OpenStudy (anonymous):

Connexus Academy.

OpenStudy (anonymous):

koo :)